OpenStudy (anonymous):

A state has a limit on the lengths of cargo trains. The train cannot exceed a length of 1,100 ft. Train A has 28 cars at 40 feet each, and train B has 34 cars at 30 feet each. Do the trains meet the requirement? a. Train A and Train B pass. c. Train A fails and Train B passes. b. Train A passes and Train B fails. d. none of these

OpenStudy (anonymous):

28x40 and 34x30 are the equations

OpenStudy (anonymous):

Train A = 28 * 40 Train B = 34 * 30

OpenStudy (anonymous):

ok so?

OpenStudy (anonymous):

28x40 is how much??

OpenStudy (anonymous):

multiply 28*40 to get the length of train A

OpenStudy (anonymous):

yes ma'am!

OpenStudy (anonymous):

a=1120

OpenStudy (anonymous):

Correct. Now, what does B equal?

OpenStudy (anonymous):

and 34x30 is equal to train B's length...

OpenStudy (anonymous):

so does train A pass or fail?

OpenStudy (anonymous):

well is it longer than 1,100 ft??

OpenStudy (anonymous):

b= 1020

OpenStudy (anonymous):

a fails?

OpenStudy (anonymous):

and is 1020 longer than 1,100 ft?

OpenStudy (anonymous):

no it's shorter

OpenStudy (anonymous):

good, yes A fails but B passes(:

OpenStudy (anonymous):

okay thx everyone

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

Train A = 28*40'. 28*40 = 1120'. 1120' is greater than 1100'. Train B = 34 * 30'. 34 * 30 = 1020' is less than 1100'.

OpenStudy (anonymous):

34*40 equales b's length